1) A project has a market beta of 1.7. The risk-free rate is 3%, and the equity premium is 5%. Your firm should undertake this project only if it returns

greater or equal to 8% greater or equal to 35% greater or equal to 8.3333% greater or equal to 11.5%

2) A zero-coupon bond has a beta of 0.3 and promises to pay $1000 next year with a probability of 95%. If the bond defaults, it will pay nothing. One -year Treasury securities are yielding 2%, and the equity premium is 5%. What is the promised rate of return on this bond? Round your answer to the nearest tenth of a percent. 8.2% 8.9% 8.0% 6.9%

3) You have analyzed the following four securities and have estimated each securityʹs beta and what you expect each security to return next year. The expected return on the market portfolio is 8%, and the relevant risk-free rate is 3%. Security Beta Expected return A 1.30 10.00% B 0.90 7.00% C 0.50 6.00% D 1.80 12.00% Based on your analysis, which of the securities is correctly priced? Security B Security D Security A Security C

4) Which of the following is not an advantage of the certainty-equivalent approach to determining the NPV of a project? It allows the decision maker to incorporate preferences for risk All of them are advantages of the certainty -equivalent approach It is easier to interpret the net present value when the certainty equivalent method is used It separates the time value of money from the risk of the project

5) Which of the following comparisons of the Capital Asset Pricing Model (CAPM) and the Arbitrage Pricing Model (APT) is (are) true? While CAPM requires a number of estimates, the APT factors and their values are known with certainty All of the are true CAPM bases the expected return on an asset on one "factor" - the market portfolio; the APT bases the expected return on a number of economic factors CAPM is based on the theory that an assetʹs returns should compensate the investor for the risk of the investment. Risk is not factored into the APT
Answer #1

1)

beta = b = 1.7

risk free rate = r = 0.03

equity premium = e = 0.05

required return = R = r + (b*e) = 0.03 +(1.7*0.05) = 0.115 = 11.5%

Thus the correct option is

greater than or equal to 11.5%

this is because the required return for the stock R = 11.5% which is the fair rate of return for the stock.